The weight of a human fetus is given by the formula W=3(t^2), where W is weight
in grams and t is the time in weeks, 0 <= t <= 39. after how many weeks does the fetus weigh 300 grams?

Answer by nerdybill(7384)

The weight of a human fetus is given by the formula W=3(t^2), where W is weight
in grams and t is the time in weeks, 0 <= t <= 39. after how many weeks does the fetus weigh 300 grams?
.
Given:
W=3(t^2)
Set W t0 300 and solve for t:
300=3(t^2)
100=t^2
10 weeks = t